write the equation in standard form, then sketch the graph of the equation: x^2=y^2-6x-4y+13=36
i dont know if you can help me with the graph or not
i hope so though
thanks a bunch

Answer by Nate(3500)

x^2 + y^2 - 6x - 4y + 13 = 36
x^2 - 6x + y^2 - 4y = 23
(x - 3)^2 + (y - 2)^2 = 23 + 9 + 4 = 36
(x - 3)^2 + (y - 2)^2 = 36
Center: (3,2)
Radius: 6